OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (2, -4) with a slope of 1/2. y =1/2 x + 5 y =1/2 x - 3 y =1/2 x - 5 y =1/2 x + 3

OpenStudy (jhannybean):

use the format for slope intercept form \[\large y=mx+b\]\[\large m= slope=\frac12\]\[\large -4=\frac12(2)+b\]solve for b (y-intercept. ) \[\large -4=1+b\]\[\large b=-5\] Now that you've found your slope and your y-intercept,plug it back into the slope intercept form.

OpenStudy (anonymous):

idk how @Jhannybean

OpenStudy (jhannybean):

\[\large y=mx+b \]\[\large m=\frac12\]\[\large b=-5\] input the values.

OpenStudy (anonymous):

Which of the following is the slope between the points (3, -2) and (8, 4)? @Jhannybean

OpenStudy (jhannybean):

Did you understand the last one?I'm not going to give you answers if that is what you're looking for.

OpenStudy (jhannybean):

To find the slope, follow the format for slope, m. \[\large m=\frac{y_{2}-y_{1}}{x_{2}-x_{1}}\]\[\large (x_{1},y_{1})=(3,-2) \ , \ (x_{2},y_{2})=(8,4)\] plug in your points and solve

OpenStudy (jhannybean):

\[\large m=\frac{4-(-2)}{8-3}=?\]

OpenStudy (anonymous):

I don't know @Jhannybean

OpenStudy (anonymous):

6/5 ?

OpenStudy (anonymous):

@Jhannybean

OpenStudy (jhannybean):

Good job.